the age of both lee and Sam is 50 years. in 5years time Sam will be 4 times the age of lee. find their ages


Answer by ikleyn(52799)

.
the age of both lee and Sam is 50 years. in 5years time Sam will be 4 times the age of lee. find their ages
~~~~~~~~~~~~~~~~~~

Let solve it MENTALLY.


Consider Lee and Sam in 5 years.


Their total age will be 60 years, and Sam will be 4 times as old as Lee.


Hence (it is obvious for a person able to calculate quickly in your mind) at that time 

Lee will be 12 years years old, while Sam will be 48 years old: 48 + 12 = 60 and 48/12 = 4.



It means that now Lee is 12-5 - 7 years old, while Sam is 48-5 = 43 years old.    ANSWER

-------------

Solved.

You can check it on your own that my answer is correct.


//////////////


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.